The initial margin requirement of an interest rate futures contract is 12% with a price of $149,841. The futures is worth $125,000 per contract. The percentage profit/loss of the investor with a short position of this futures will be _________ if the futures price becomes $145,000.

Multiple Choice

A.37.92% loss

B. 37.92% profit

C. 26.92% profit

D. 26.92% loss
Answer #1

Initial Margin = 0.12(149,841) = $17,980.92

Profit = (149,841 - 145,000) = $4,841

Profit % = 4,841/17,980.92

Profit % = 26.92%